OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20

OpenStudy (anonymous):

|x + 2| + 16 = 14 |x + 2| = 14-16 |x + 2| = -2 Since the absolute value of any number is nonnegative, this means that the last equation is not possible So there are no solutions.

OpenStudy (solomonzelman):

\(\large\color{black}{|2x + 6| − 4 = 20 }\) \(\large\color{black}{|2x + 6| − 4 \color{magenta}{+4}= 20\color{magenta}{+4} }\) \(\large\color{black}{|2x + 6| = 24 }\) \(\large\color{black}{|2(x +3)| = 24 }\) \(\large\color{black}{2|x +3| = 24 }\) \(\large\color{black}{2|x +3| \color{magenta}{\div 2} = 24 \color{magenta}{\div 2} }\) \(\large\color{black}{|x +3| = 12 }\) \(\large\color{black}{x +3 = 12 ~~~~~~or~~~~~x+3=-12 }\) \(\large\color{black}{x = 9 ~~~~~~~~~~~~~~~~~~~~~~~~x=-15 }\)

OpenStudy (anonymous):

so my answer is no solution?

OpenStudy (solomonzelman):

why?? I gave you the solutions with steps....